# Prompt and Parameter Co-Optimization for Large Language Models

### v1

@@ Abstract
Prompt optimization and fine-tuning are two major approaches to improve the performance of Large Language Models (LLMs). They enhance the capabilities of LLMs from complementary perspectives: the former through explicit natural language, and the latter through implicit parameter updates. However, prior work has typically studied them in isolation, leaving their synergistic potential largely underexplored. To bridge this gap, in this paper, we introduce MetaTuner, a novel framework that jointly integrates prompt optimization and fine-tuning for LLM training. Specifically, we introduce two neural networks to generate prompts and parameters, respectively, while allowing them to share a common bottom encoding layer to enable knowledge sharing. By the guidance of the final supervised signals, our framework is optimized to discover the optimal combinations between the prompts and parameters. Given that prompt learning involves discrete optimization while fine-tuning operates in a continuous parameter space, we design a supervised regularization loss to train our framework effectively. Extensive experiments across diverse benchmarks show that our method consistently outperforms the baselines.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success across a wide range of domains, including e-commerce (ref:shoppingmmlu, ; ref:ecellm, ; ref:ecomgpt, ; ref:ecommerce, ), education (ref:edu1, ; ref:edu2, ; ref:edu3, ), and social science (ref:social1, ; ref:social2, ; ref:social3, ; ref:social4, ). Unlike traditional machine learning models, whose capabilities primarily stem from parameter tuning, LLMs can enhance their performance through two complementary strategies: prompt optimization and fine-tuning.
@@
Previously, prompt optimization and fine-tuning are often treated as independent research lines. For example, in the line of prompt optimization, current methods (ref:hmaw, ; ref:grad_sum, ; ref:llm_dual_phase_po, ; ref:rlprompt_improve, ) mainly focus on discovering or learning appropriate input context that can elicit the desired behavior from frozen language models. One group of works (e.g., OPRO (ref:opro, )) focuses on optimizing prompts with LLMs, whereas another group (e.g., RLPrompt (ref:rlprompt, ), BPO (ref:bpo, )) trains an auxiliary model for prompt generation. In the line of fine-tuning, existing approaches concentrate on better training algorithms to adjust the model parameters. Supervised Fine-Tuning (SFT) proposes to optimize the model with cross-entropy loss, while more recently, RL-based approaches such as RLHF (ref:rlhf, ) and DPO (ref:dpo, ) are developed to optimize the model based on human feedback signals.
@@
While the above models have shown promising results in various domains, both prompt optimization and fine-tuning have their own limitations. Specifically, prompt optimization typically focuses on refining general instructions within prompts to elicit the capabilities of LLMs. However, it often falls short in adapting the pretrained parameters to the complex patterns present in large-scale, task-specific datasets, especially when the knowledge encoded in the prompts and the model parameters may conflict. Fine-tuning, on the other hand, usually leverages manually crafted prompts as inputs during model training. However, the choice of input prompts can significantly affect fine-tuning outcomes, and existing methods offer no guarantee that these prompts are optimal for maximizing downstream performance. To further validate these statements, we conduct a series of preliminary experiments, and the corresponding results are presented in Figure 1(a). As illustrated in the left subfigure, by training the LLM parameters based on the datasets of MATH (ref:math, ) and HotpotQA (ref:hotpotqa, ), the fine-tuning methods substantially outperform the prompt optimization strategies on average. However, the effectiveness of fine-tuning is highly sensitive to the choice of prompts. As shown in the right subfigure, when sub-optimal prompts are used for training, the performance degrades and can even fall below that of prompt optimization methods. Motivated by the above mechanistic and empirical analyses, a natural question arises: Can we design a unified framework to combine prompt optimization and fine-tuning to mitigate the weaknesses of each other?
@@ 2 Problem Formulation
In LLMs, we typically work with an input-output pair $(x,y)$, where $x$ represents a question such as “What is the capital of France?”, and $y$ corresponds to the expected answer like “Paris”. Unlike traditional machine learning models, LLMs often require an additional prompt $p$ to rephrase the input $x$ in a way that better guides the model to generate the desired output $y$. In the example above, $p$ could be “Answer the following question briefly:”. Letting the language model be denoted as $\mathcal{M}$, the overall process can be formulated as:
@@
Previous studies have employed various strategies (ref:rlhf, ; ref:dpo, ; ref:kto, ; ref:opro, ; ref:bpo, ; ref:rlprompt, ; ref:dora, ; ref:pissa, ) to instantiate objectives (2) and (3), respectively. Here, we take a different perspective by formulating them within a unified framework, laying the foundation for letting prompt optimization and fine-tuning mutually enhance each other. Specifically, we regard the prompt as a “special parameter”, then objectives (2) and (3) can be unified by the following loss function:
@@
Compared with previous studies, this loss function considers prompts and model parameters as two complementary dimensions that jointly influence the prediction from $x_{i}$ to $y_{i}$. By optimizing them within a unified objective, we are actually finding the optimal combinations between the prompts and parameters to better fit specific tasks. The loss function can help to regularize any sub-optimal choices in either the prompt or parameter space. Readers may find that the above objective can also be applied to soft-prompting strategies (ref:p_tuning, ; ref:p_tuning_v2, ; ref:prefix_tuning, ; ref:google_soft_prompt, ; ref:prompt_tuning, ). However, in this paper, we focus on generating explicit prompts, because the natural language prompts are more explainable, transparent, and human-friendly.
@@ 3.3 Optimization with Supervised Regularization
To address the above non-differentiable problem, we also explored strategies such as Gumbel-Softmax (ref:gumbel_softmax, ), but these methods generally produced inferior performance. We speculate that our supervised regularizer is more effective because it provides direct supervision on the optimal prompts, avoiding intermediate softening steps or approximations.
@@ 3.4 Framework Specification
Specification of $\mathcal{G}$. In our framework, an LLM such as Qwen2.5-7B (ref:qwen2.5, ) is selected as the implementation of $\mathcal{G}$. In specific, we use the first $k$ decoder layers as the meta encoder (i.e., $\phi_{s}$), while the subsequent layers serve as the prompt decoder (i.e., $\phi_{p}$). We denote the hidden states by $h_{i}\in\mathbb{R}^{l\times h_{\mathcal{G}}}$, where $l$ represents the input sequence length and $h_{\mathcal{G}}$ corresponds to the hidden size.
Specification of $\mathcal{F}$. For training efficiency, we adopt LoRA (ref:lora, ) to fine-tune the downstream LLMs, with $\mathcal{F}$ responsible for generating the LoRA parameters. Specifically, let $\Delta W=\theta_{i}^{b}\cdot\theta_{i}^{a}$ denote the update matrix, where $\theta_{i}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times r_{\mathcal{M}}}$ and $\theta_{i}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times k_{\mathcal{M}}}$. $\mathcal{F}$ first share the same $\phi_{s}$ with $\mathcal{G}$ to produce $h_{i}$. After that, we derive ${\theta}^{b}_{i}$ and ${\theta}^{a}_{i}$ from the hidden states $h_{i}$ as follows:
@@ 4.1 Experimental Setup
Datasets and metrics. We evaluate MetaTuner on four datasets: MATH (ref:math, ) and GSM8K (ref:gsm8k, ) for mathematical reasoning, HotpotQA (ref:hotpotqa, ) and CosmosQA (ref:cosmosqa, ) for multi-hop and commonsense question answering. Each dataset is divided into training, development, and test sets, with details of the splits provided in the Appendix D.2. We report the exact match (EM) accuracy for MATH, GSM8K, and CosmosQA, while for HotpotQA, we adopt the F1 accuracy following the implementation of (ref:copper, ; ref:retroformer, ).
@@
Baselines. We compare MetaTuner with a series of state-of-the-art approaches, which can be classified into four categories: (1) Vanilla Methods: Qwen2.5 (ref:qwen2.5, ) and GPT-4o (ref:gpt4o, ) are employed as essential baselines to directly answer queries under the zero-shot setting. (2) Prompt Optimization Methods: RLPrompt (ref:rlprompt, ), BPO (ref:bpo, ), OPRO (ref:opro, ), and CFPO (ref:cfpo, ). For RLPrompt and BPO, we use Qwen2.5 as the base prompt generator, while for LLM-based methods such as OPRO and CFPO, we use GPT-4o as the prompt optimizer. (3) Fine-Tuning Methods: SFT, PPO (ref:ppo, ), DPO (ref:dpo, ), and KTO (ref:kto, ). For RL-based methods (PPO, DPO, and KTO), we first perform SFT as a warm-up stage and then construct alignment data following SPIN (ref:spin, ). (4) Hybrid Methods: BetterTogether (ref:better_together, ) is a hybrid method that iteratively optimizes the prompts and weights. However, it does not incorporate the parameter generation network and knowledge-sharing mechanism. As mentioned in the Section 3.2, we optimize our model with two strategies: In MetaTuner-I, we adopt a sequential optimization strategy, where the two terms in Equation 8 are optimized alternatively. In MetaTuner-J, we optimize Equation 8 in a unified manner.
@@
Implementation details. We use Qwen2.5-7B and Qwen2.5-3B as the prompt generator $\mathcal{G}$ and the downstream actor model $\mathcal{M}$. Prior to joint training, we initiate an SFT stage to warm up $\mathcal{G}$ and $\mathcal{M}$ respectively. The training is carried out through Llama-Factory (ref:llamafactory, ), with the epoch set to $1$, batch size set to $64$, and the learning rate tuned in {1ee-3, 5ee-4, 1ee-4, 5ee-5}. Details of the stage can be found in the Appendix D.2. We then search the depth $k$ of $\phi_{s}$ in $\{K/4,K/2,3K/4,K\}$ to determine the model structure, where $K$ denotes the number of decoder layers in $\mathcal{G}$. In the training process, we search $r_{\mathcal{M}}$ in {4, 8, 16, 32}, $\alpha$ in {0.1, 0.5, 0.9}, the temperature $t$ for online rollouts in {0, 0.3, 0.5, 0.7, 0.9}, and the learning rate in {1ee-5, 5ee-6, 1ee-6, 5ee-7}. Additionally, we introduce a hyper-parameter $\lambda$ to scale the generated LoRA before adding it to the original model weights, and search $\lambda$ over {0.01, 0.05, 0.1, 0.5}. During the inference stage, we employ greedy decoding to ensure reproducibility. All the experiments are conducted on eight NVIDIA A100-80G GPUs.
@@ 4.5 Effectiveness of the Supervised Regularization Loss
In our framework, a key design for addressing the mixed discrete–continuous optimization problem is the incorporation of the supervised regularizer in the objective (8). To evaluate its effectiveness, we conduct comprehensive experiments based on four research questions: (1) Is our method superior to other well-known differentiable strategies such as Gumbel-Softmax? (2) In selecting the dataset $D_{2}$, should it be derived from the optimized model itself or from a stronger model such as GPT-4o? (3) How does the number of rollout samples affect performance? (4) How does the update frequency of $\phi_{p}$ influence the final performance? In the following, we report and analyze the experiment results.
@@
Comparing with other differentiable strategies. In this section, we compare our method with Gumbel-Softmax, where the categorical prompt token sampling is replaced with a Gumbel-Softmax distribution to enable differentiable sampling. As shown in Figure 3, our method significantly outperforms Gumbel-Softmax. The main reason is that Gumbel-Softmax relies on continuous relaxation to approximate the discrete sampling. Although this ensures differentiability in theory, such softening operations introduce gradient bias, so that the generated prompts may deviate semantically from the truly optimal discrete solutions. In contrast, our methods directly optimize over the discrete action space using task rewards, thereby avoiding the potential approximation errors.
@@
Comparison between different generation methods of $D_{2}$. We compare the performances by generating $D_{2}$ using GPT-4o or the optimized model itself. As shown in Figure 4, while the GPT-4o method improves training efficiency, our model achieves better performance. The key reasons are threefold: first, our model ensures distribution alignment by generating data directly from the current policy, thereby reducing distribution shift between training and inference; second, it provides more targeted feedback, as rewards are based on the model’s own outputs and directly address its weaknesses, unlike GPT-4o samples that may overlook error patterns; third, our model enables stronger exploration, allowing the policy to continuously discover new prompts and iteratively refine itself through task feedback, whereas the GPT-4o data remain static and limit long-term improvement.

### v2

# arXiv:2509.24245v2[cs.CL] 02 Mar 2026

@@ Abstract
Prompt optimization and fine-tuning are two major approaches to improve the performance of Large Language Models (LLMs). They enhance the capabilities of LLMs from complementary perspectives: the former through explicit natural language, and the latter through implicit parameter updates. However, prior work has typically studied them in isolation, leaving their synergistic potential largely underexplored. To bridge this gap, in this paper, we introduce MetaTuner, a novel framework that jointly integrates prompt optimization and fine-tuning for LLM training. Specifically, we introduce two neural networks to generate prompts and parameters, respectively, while allowing them to share a common bottom encoding layer to enable knowledge sharing. By the guidance of the final supervised signals, our framework is optimized to discover the optimal combinations between the prompts and parameters. Given that prompt learning involves discrete optimization while fine-tuning operates in a continuous parameter space, we design a supervised regularization loss to train our framework effectively. Extensive experiments across diverse benchmarks show that our method consistently outperforms the baselines. To benefit the research community, we have released our project at https://github.com/BoXiaohe/MetaTuner.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success across a wide range of domains, including e-commerce (Jin et al., 2024; Peng et al., 2024; Li et al., 2024; Ren et al., 2024), education (Wang et al., 2024a; Ghimire et al., 2024; Chu et al., 2025), and social science (Wang et al., 2025; Mou et al., 2025; Tang et al., 2024; Piao et al., 2025). Unlike traditional machine learning models, whose capabilities primarily stem from parameter tuning, LLMs can enhance their performance through two complementary strategies: prompt optimization and fine-tuning.
@@
Previously, prompt optimization and fine-tuning are often treated as independent research lines. For example, in the line of prompt optimization, current methods (Liu et al., 2025b; Austin and Chartock, 2024; Yang et al., 2024b; Choi et al., 2024) mainly focus on discovering or learning appropriate input context that can elicit the desired behavior from frozen language models. One group of works (e.g., OPRO (Yang et al., 2024a)) focuses on optimizing prompts with LLMs, whereas another group (e.g., RLPrompt (Deng et al., 2022), BPO (Cheng et al., 2024)) trains an auxiliary model for prompt generation. In the line of fine-tuning, existing approaches concentrate on better training algorithms to adjust the model parameters. Supervised Fine-Tuning (SFT) proposes to optimize the model with cross-entropy loss, while more recently, RL-based approaches such as RLHF (Li et al., 2023b) and DPO (Rafailov et al., 2023) are developed to optimize the model based on human feedback signals.
@@
While the above models have shown promising results in various domains, both prompt optimization and fine-tuning have their own limitations. Specifically, prompt optimization typically focuses on refining general instructions within prompts to elicit the capabilities of LLMs. However, it often falls short in adapting the pretrained parameters to the complex patterns present in large-scale, task-specific datasets, especially when the knowledge encoded in the prompts and the model parameters may conflict. Fine-tuning, on the other hand, usually leverages manually crafted prompts as inputs during model training. However, the choice of input prompts can significantly affect fine-tuning outcomes, and existing methods offer no guarantee that these prompts are optimal for maximizing downstream performance. To further validate these statements, we conduct a series of preliminary experiments, and the corresponding results are presented in Figure 1(a). As illustrated in the left subfigure, by training the LLM parameters based on the datasets of MATH (Hendrycks et al., 2021) and HotpotQA (Yang et al., 2018), the fine-tuning methods substantially outperform the prompt optimization strategies on average. However, the effectiveness of fine-tuning is highly sensitive to the choice of prompts. As shown in the right subfigure, when sub-optimal prompts are used for training, the performance degrades and can even fall below that of prompt optimization methods. Motivated by the above mechanistic and empirical analyses, a natural question arises: Can we design a unified framework to combine prompt optimization and fine-tuning to mitigate the weaknesses of each other?
@@ 2 Problem Formulation
In the post-training of LLMs, we typically work with an input-output pair $(x,y)$, where $x$ represents a question such as “What is the capital of France?”, and $y$ corresponds to the expected answer like “Paris”. Unlike traditional machine learning models, LLMs often require an additional prompt $p$ to rephrase the input $x$ in a way that better guides the model to generate the desired output $y$. In the example above, $p$ could be “Answer the following question briefly:”. Let $\mathcal{M}$ be the language model, the overall process can be written as:
@@
Previous studies have employed various strategies (Li et al., 2023b; Rafailov et al., 2023; Ethayarajh et al., 2024; Yang et al., 2024a; Cheng et al., 2024; Deng et al., 2022; Liu et al., 2024a; Meng et al., 2024) to instantiate objectives (2) and (3), respectively. Here, we take a different perspective by formulating them within a unified framework, laying the foundation for letting prompt optimization and fine-tuning mutually enhance each other. Specifically, we regard the prompt as a “special parameter”, then objectives (2) and (3) can be unified by the following loss function:
@@
Compared with previous studies, this loss function considers prompts and model parameters as two complementary dimensions that jointly influence the prediction from $x_{i}$ to $y_{i}$. By optimizing them within a unified objective, we are actually finding the optimal combinations between the prompts and parameters to better fit specific tasks. The loss function can help to regularize any sub-optimal choices in either the prompt or parameter space. Readers may find that the above objective can also be applied to soft-prompting strategies (Liu et al., 2024b; 2022; Li and Liang, 2021; Lester et al., 2022; 2021). However, in this paper, we focus on generating explicit prompts, because the natural language prompts are more explainable, transparent, and human-friendly.
@@ 3.3 Optimization with Supervised Regularization
To address the above non-differentiable problem, we also explored strategies such as Gumbel-Softmax (Jang et al., 2017), but these methods generally produced inferior performance. We speculate that our supervised regularizer is more effective because it provides direct supervision on the optimal prompts, avoiding intermediate softening steps or approximations.
@@ 3.4 Framework Specification
Specification of $\mathcal{G}$. In our framework, an LLM such as Qwen2.5-7B (Team, 2024) is selected as the implementation of $\mathcal{G}$. In specific, we use the first $k$ decoder layers as the meta encoder (i.e., $\phi_{s}$), while the subsequent layers serve as the prompt decoder (i.e., $\phi_{p}$). We denote the hidden states by $h_{i}\in\mathbb{R}^{l\times h_{\mathcal{G}}}$, where $l$ represents the input sequence length and $h_{\mathcal{G}}$ corresponds to the hidden size.
Specification of $\mathcal{F}$. For training efficiency, we adopt LoRA (Hu et al., 2022) to fine-tune the downstream LLMs, with $\mathcal{F}$ responsible for generating the LoRA parameters. Specifically, let $\Delta W=\theta_{i}^{b}\cdot\theta_{i}^{a}$ denote the update matrix, where $\theta_{i}^{b}\in\mathbb{R}^{d_{\mathcal{M}}\times r_{\mathcal{M}}}$ and $\theta_{i}^{a}\in\mathbb{R}^{r_{\mathcal{M}}\times k_{\mathcal{M}}}$. $\mathcal{F}$ first share the same $\phi_{s}$ with $\mathcal{G}$ to produce $h_{i}$. After that, we derive ${\theta}^{b}_{i}$ and ${\theta}^{a}_{i}$ from the hidden states $h_{i}$ as follows:
@@ 4 Experiments
  • RQ6: How does MetaTuner perform compared to the soft co-optimization approach?
@@ 4.1 Experimental Setup
Datasets and metrics. We evaluate MetaTuner on four datasets: MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) for mathematical reasoning, HotpotQA (Yang et al., 2018) and CosmosQA (Li et al., 2023a) for multi-hop and commonsense question answering. Each dataset is divided into training, development, and test sets, with details of the splits provided in the Appendix D.2. We report the exact match (EM) accuracy for MATH, GSM8K, and CosmosQA, while for HotpotQA, we adopt the F1 accuracy following the implementation of (Bo et al., 2024; Yao et al., 2024).
@@
Baselines. We compare MetaTuner with a series of state-of-the-art approaches, including: (1) Vanilla Methods: Qwen2.5 (Team, 2024) is employed as an essential baseline to directly answer queries under the zero-shot setting. (2) Prompt Optimization Methods: RLPrompt (Deng et al., 2022), BPO (Cheng et al., 2024), OPRO (Yang et al., 2024a), and CFPO (Liu et al., 2025a). For RLPrompt and BPO, we use Qwen2.5 as the base prompt generator, while for LLM-based methods such as OPRO and CFPO, we use state-of-the-art proprietary LLMs as the prompt optimizer. (3) Fine-Tuning Methods: SFT, PPO (Ouyang et al., 2022), DPO (Rafailov et al., 2023), and KTO (Ethayarajh et al., 2024). For RL-based methods (PPO, DPO, and KTO), we first perform SFT as a warm-up stage and then construct alignment data following SPIN (Chen et al., 2024). (4) Hybrid Methods: BetterTogether (Soylu et al., 2024) is a hybrid method that iteratively optimizes the prompts and weights. However, it does not incorporate the parameter generation network and knowledge-sharing mechanism. As mentioned in the Section 3.2, we optimize our model with two strategies: In MetaTuner-I, we adopt a sequential optimization strategy, where the two terms in Equation 8 are optimized alternatively. In MetaTuner-J, we optimize Equation 8 in a unified manner.
@@
Implementation details. We use Qwen2.5-7B and Qwen2.5-3B as the prompt generator $\mathcal{G}$ and the downstream actor model $\mathcal{M}$. Prior to joint training, we initiate an SFT stage to warm up $\mathcal{G}$ and $\mathcal{M}$ respectively. The training is carried out through Llama-Factory (Zheng et al., 2024), with the epoch set to $1$, batch size set to $64$, and the learning rate tuned in {1ee-3, 5ee-4, 1ee-4, 5ee-5}. Details of the stage can be found in the Appendix D.2. We then search the depth $k$ of $\phi_{s}$ in $\{K/4,K/2,3K/4,K\}$ to determine the model structure, where $K$ denotes the number of decoder layers in $\mathcal{G}$. In the training process, we search $r_{\mathcal{M}}$ in {4, 8, 16, 32}, $\alpha$ in {0.1, 0.5, 0.9}, the temperature $t$ for online rollouts in {0, 0.3, 0.5, 0.7, 0.9}, and the learning rate in {1ee-5, 5ee-6, 1ee-6, 5ee-7}. Additionally, we introduce a hyper-parameter $\lambda$ to scale the generated LoRA before adding it to the original model weights, and search $\lambda$ over {0.01, 0.05, 0.1, 0.5}. During the inference stage, we employ greedy decoding to ensure reproducibility. All the experiments are conducted on eight 80G GPUs.
@@ 4.5 Effectiveness of the Supervised Regularization Loss
In our framework, a key design for addressing the mixed discrete–continuous optimization problem is the incorporation of the supervised regularizer in the objective (8). To evaluate its effectiveness, we conduct comprehensive experiments based on four research questions: (1) Is our method superior to other well-known differentiable strategies such as Gumbel-Softmax? (2) In selecting the dataset $D_{2}$, should it be derived from the optimized model itself or from a stronger model? (3) How does the number of rollout samples affect performance? (4) How does the update frequency of $\phi_{p}$ influence the final performance? In the following, we report and analyze the experiment results.
@@
Comparing with other differentiable strategies. We compare our method with Gumbel-Softmax, where the categorical prompt token sampling is replaced with a Gumbel-Softmax distribution to enable differentiable sampling. As shown in Figure 3, our method significantly outperforms Gumbel-Softmax. The main reason is that Gumbel-Softmax relies on continuous relaxation to approximate the discrete sampling. Although it ensures differentiability in theory, such softening operations introduce gradient bias, so that the generated prompts may deviate semantically from the truly optimal discrete solutions. In contrast, our methods directly optimize over the discrete action space, avoiding potential approximation errors.
@@
Comparison between different generation methods of $D_{2}$. We compare the performances by generating $D_{2}$ using the proprietary expert LLM or the optimized model itself. As shown in Figure 4, while the expert method improves training efficiency, our model achieves better performance. The key reasons are threefold: first, our model ensures distribution alignment by generating data directly from the current policy, thereby reducing distribution shift between training and inference; second, it provides more targeted feedback, as rewards are based on the model’s own outputs and directly address its weaknesses, unlike expert samples that may overlook error patterns; third, our model enables stronger exploration, allowing the policy to continuously discover new prompts and iteratively refine itself through task feedback, whereas the expert data remain static and limit long-term improvement.
@@ 4.6 Generalization Analysis
4.7 Comparison with Soft Co-Optimization Method
We compare MetaTuner with soft co-optimization methods and report the results in Figure 8(a). We employ prefix tuning (Li and Liang, 2021) for soft prompt optimization, while using LoRA as the fine-tuning approach, and fine-tune the model with SFT. From the results, we can find that the optimization of discrete prompts is more stable. Existing work (Bhardwaj et al., 2022; Razdaibiedina et al., 2023) has shown that soft prompt tuning can suffer from “representation collapse”, where prompts for different inputs converge to nearly identical vectors. In MetaTuner, we optimize discrete prompts, i.e., prompts must take the form of readable natural language, with each token subject to clear semantic constraints. This forces the optimization process to search over structured, human-interpretable instructions rather than arbitrary continuous vectors, thereby reducing the risk of overfitting.
(a) Comparing MetaTuner with Soft Co-Optimization method.